OpenStudy (anonymous):

What is the sum of the geometric sequence 4, 16, 64, … if there are 8 terms? 21,845 43,690 65,535 87,380

OpenStudy (texaschic101):

the numbers are multiplied by 4..... start with 4 (1st term) 4 (4) = 16 (2nd term) 16 (4) = 64 (3rd term) 64 (4) = 256 (4th term) 256 (4) = 1024 (5th term) 1024 (4) = 4096 (6th term) 4096 (4) = 16384 (7th term) 16384 (4) = 65536 (8th term) Now you want the total, so just add... 4 + 16 + 64 + 256 + 1024 + 4096 + 16384 + 65536 = 87380 There is probably an easier way to do this....I just don't know it. I couldn't have done this if there were 50 terms...lol

OpenStudy (kropot72):

You can use the formula \[S _{n}=\frac{a(r ^{n}-1)}{r-1}\]

OpenStudy (anonymous):

Thank you!

OpenStudy (kropot72):

You're welcome :)